# Boosting Robustness by Clipping Gradients in Distributed Learning

### v1

@@ Appendix C Proofs for Results in Section 5
Consider an arbitrary $t\in[T]$. Let $(x_{1},\ldots,\,x_{n})\coloneqq\mathbf{ARC}\left({{g}^{\left({1}\right)}_{t}},\ldots,~{}{{g}^{\left({n}\right)}_{t}}\right)$. As $n>2f$ and $\left\lvert{\mathcal{H}}\right\rvert=n-f$, the clipping threshold $C$ used in ARC (see Algorithm 2) is bounded by $\max_{i\in\mathcal{H}}\left\lVert{{{g}^{\left({i}\right)}_{t}}}\right\rVert$. Therefore,
@@
Since we assume (without loss of generality) that, for any set of $n$ vectors $z_{1},\ldots,z_{n}\in\mathbb{R}^{d}$, $\left\lVert{\mathbf{F}\left(z_{1},\ldots,z_{n}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{z_{i}}\right\rVert$ (see the discussion in Section 5.1), from (16) we obtain that
@@
We are now ready to present our proof of Theorem 5.1, recalled below.
See 5.1
Proof.
For simplicity we write $\mathcal{L}_{\mathcal{H}}$ as $\mathcal{L}$ throughout the proof.
Consider an arbitrary $t\in[T]$. Due to $L$-Lipschitz smoothness of $\mathcal{L}({{\theta}_{)}}$, we obtain that
@@
$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq\frac{2\Delta_{o}}{\gamma T}+\frac{3\kappa}{T}~{}\frac{1}{\left\lvert{\mathcal{H}}\right\rvert}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})-\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}.$ (18)
@@
C.3 Proof of Corollary 5.2
See 5.2
Proof.
Since $T\geq\frac{\Delta_{o}L}{\kappa G^{2}}$, $\gamma=\frac{\Delta_{o}}{\kappa G^{2}T}\leq\frac{1}{L}$. Therefore, from Theorem 5.1 we obtain that
$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq 5\left(G^{2}+B^{2}\rho^{2}\right).$ (22)
@@
C.4 Proof of Theorem 5.4
Recall that we denote
$\mathsf{BP}\coloneqq\frac{1}{2+B^{2}},\quad\varepsilon_{o}\left(f/n\right)\coloneqq\frac{1}{4}\cdot\frac{G^{2}(f/n)}{1-(2+B^{2})(f/n)},\quad\text{ and }$
@@
We recall the theorem below for convenience.
See 5.4
Proof.
@@
First, we consider $\frac{f}{n}=\frac{1-\xi}{2+B^{2}}$. Note that, due to Lemma 3.3, $\mathbf{F}$ is $(f,\kappa)$-robust with
$\kappa\leq\frac{16f}{n-f}\left(1+\frac{1-\xi}{B^{2}}\right)^{2}\leq\frac{16(1-\xi)}{1+B^{2}}\left(1+\frac{1-\xi}{B^{2}}\right)^{2}\leq\kappa_{o}(1-\xi),$
@@
To summarize the above, we have
@@
Therefore, since $\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\zeta$ for all $i\in\mathcal{H}$ under Assumption 5.3, in Theorem 5.1 we always satisfy the condition: $\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\exp\left(-\frac{\Delta_{o}}{\kappa G^{2}}L\right)\rho$, for the defined value of $\rho$ Thus, by Corollary 5.2 and (24) we obtain that
@@
Substituting from above in (25) concludes the proof for the first part of the theorem.
The second part of the theorem follows immediately from the first inequality in (25), using the fact that $\kappa_{o}(1-\xi)\leq\kappa_{o}$. This concludes the proof. ∎
C.5 Proof of Corollary 5.5
See 5.5
Proof.
The proof follows immediately from upon using $\xi\leq\frac{1}{\Psi(G,B,\rho)}\upsilon$ in Part 1 of Theorem 5.4. ∎
@@ D.1 Datasets and Heterogeneity
In our experiments, we consider three standard image classification datasets, namely MNIST [11], Fashion-MNIST [42], and CIFAR-10 [31]. To simulate data heterogeneity in our experiments, we make the honest workers sample from the datasets using a Dirichlet [23] distribution of parameter $\alpha$, as done in [24, 3, 13]. The smaller the $\alpha$, the more heterogeneous the setting. In our empirical evaluation, we set $\alpha\in\{0.1,0.5,1\}$ on (Fashion-)MNIST and $\alpha\in\{0.05,0.075,0.1,0.2,0.5\}$ on CIFAR-10 (refer to Figure 5). Furthermore, on MNIST and Fashion-MNIST, we also consider an extreme heterogeneity setting where the datapoints are sorted by increasing labels (0 to 9) and sequentially split equally among the honest workers.
@@ D.2 Algorithm, Distributed System, ML Models, and Hyperparameters
We perform our experiments using Robust-DSGD (see Algorithm 3), which is known to be an order-optimal robust variant of Robust-DGD [3]. On MNIST and Fashion-MNIST, we execute Robust-DSGD in a distributed system composed of $n-f=10$ honest workers, and $f\in\{1,2,3,4,5,7,9\}$ adversarial workers. Furthermore, we train a convolutional neural network (CNN) of 431,080 parameters with batch size $b=25$, $T=1000$, $\gamma=0.1$, and momentum parameter $\beta=0.9$. Moreover, the negative log likelihood (NLL) loss function is used, along with an $\ell_{2}$-regularization of $10^{-4}$. On CIFAR-10, we execute Robust-DSGD using a CNN of 1,310,922 parameters, in a distributed system comprising $n-f=16$ honest workers and $f=1$ adversarial worker. We set $b=50$, $T=2000$, $\beta=0.9$, and $\gamma=0.05$ decaying once at step 1500. Finally, we use the NLL loss function with an $\ell_{2}$ regularization of $10^{-2}$.
Algorithm 3 Robust Stochastic Distributed Gradient Descent (Robust-DSGD)
@@
In order to present the architectures of the ML models used in our experiments, we adopt the following compact terminology introduced in [3].
@@ D.3 Byzantine Attacks
In our experiments, the adversarial workers execute five state-of-the-art adversarial attacks from the Byzantine ML literature, namely sign-flipping (SF) [2], label-flipping (LF) [2], mimic [27], fall of empires (FOE) [43], and a little is enough (ALIE) [6].
@@
  • Mimic: the adversarial workers mimic a certain honest worker by sending its gradient to the server. In order to determine the optimal honest worker for the adversarial workers to mimic, we use the heuristic in [27].
@@ Appendix F Static Clipping vs ARC
(a) Sign-flipping (SF) [2]
(b) Label-flipping (LF) [2]
Figure 15: Impact of pre-aggregation clipping, specifically static clipping with $C=2$ and ARC, on robust-DGD with aggregation scheme $\mathbf{CWTM}\circ\mathbf{NNM}$, under 2 adversarial attacks: (a) SF and (b) LF. We set $n=13$, $f=3$, and consider the MNIST [11] dataset distributed amongst $n-f=10$ honest workers with extreme heterogeneity.
@@ F.1 Limitations of Static Clipping
We perform our experiments using Robust-DSGD (see Algorithm 3) on (Fashion-)MNIST and CIFAR-10. We consider three different levels of heterogeneity, that we call moderate (corresponding to sampling from a Dirichlet distribution of parameter $\alpha=1$), high ($\alpha=0,1$), and extreme (as explained in Appendix D). On MNIST and Fashion-MNIST, we execute Robust-DSGD in a distributed system composed of $n=15$ workers, among which $f\in\{3,4\}$ are Byzantine. Furthermore, we train a convolutional neural network of 431,080 parameters with batch size $b=25$, $T=1000$, $\gamma=0.1$, and momentum parameter $\beta=0.9$. Moreover, the negative log likelihood loss function is used, along with an $\ell_{2}$-regularization of $10^{-4}$. On CIFAR-10, we execute Robust-DSGD on ResNet-18 [21], in a distributed system comprising $n=9$ workers among which $f\in\{1,2\}$ are Byzantine. We set $b=128$, $T=2000$, $\beta=0.9$, and $\gamma=0.1$ decaying once by 10$\times$ at step 1500. Finally, we use the cross-entropy loss function with an $\ell_{2}$ regularization of $5\times 10^{-4}$.
@@
Unreliability due to the attack. First, the efficacy of static clipping is significantly influenced by the type of Byzantine attack executed during the learning process. To illustrate, while $C=2$ exhibits the best performance under the SF attack [2], depicted in Figure 15, it leads to a complete collapse of learning under LF [2]. Consequently, identifying a static clipping threshold $C$ that consistently performs well in practice against any attack is difficult (if not impossible). This highlights the fragility of static clipping, as the unpredictable nature of the Byzantine attack, a parameter that cannot be a priori known, can significantly degrade the performance of the chosen static clipping approach.
@@ F.2 Empirical Benefits of ARC
ARC exhibits robust performance across diverse Byzantine scenarios, encompassing variations in both the type of Byzantine attack and the number of Byzantine workers $f$. As depicted in Figures 15 and 16, ARC consistently yields robust performance, regardless of the Byzantine attack. In extreme heterogeneity with $f=3$, ARC maintains an accuracy of 75%, while $C=2$ demonstrates a subpar worst-case accuracy of 25% (also refer to the right plot of Figure 15). Furthermore, despite the increase in the number of Byzantine workers to $f=4$, ARC remains the top-performing clipping approach among all considered strategies.

### v3

@@ Appendix C Proofs of the Results in Section 5
Consider an arbitrary $t\in[T]$. Let $(x_{1},\ldots,\,x_{n})\coloneqq\mathbf{ARC}\left({{g}^{\left({1}\right)}_{t}},\ldots,~{}{{g}^{\left({n}\right)}_{t}}\right)$. As $n>2f$ and $\left\lvert{\mathcal{H}}\right\rvert=n-f$, the clipping threshold $C$ used in ARC (see Algorithm 2) is bounded by $\max_{i\in\mathcal{H}}\left\lVert{{{g}^{\left({i}\right)}_{t}}}\right\rVert$ (see (20) in Lemma 5.1). Therefore,
@@
Since for any set of $n$ vectors $z_{1},\ldots,z_{n}\in\mathbb{R}^{d}$, $\left\lVert{\mathbf{F}\left(z_{1},\ldots,z_{n}\right)}\right\rVert\leq\max_{i\in[n]}\left\lVert{z_{i}}\right\rVert$, from (23) we obtain that
@@
We are now ready to present our proof of Lemma C.2.
Proof of Lemma C.2.
For simplicity, we write $\mathcal{L}_{\mathcal{H}}$ as $\mathcal{L}$ throughout the proof.
Consider an arbitrary $t\in[T]$. Due to $L$-Lipschitz smoothness of $\mathcal{L}({{\theta}})$, we obtain that
@@
$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq\frac{2\Delta_{o}}{\gamma T}+\frac{3\kappa}{T}~{}\sum_{t=1}^{T}\frac{1}{\left\lvert{\mathcal{H}}\right\rvert}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{t}})-\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}.$ (25)
@@
We now prove Theorem C.4, stated below, which immediately implies Theorem 5.2. Specifically, Theorem 5.2 follows immediately from Part 1 of Theorem C.4, upon substituting $\xi\leq\frac{1}{\Psi(G,B,\rho)}\upsilon$.
Theorem C.4.
Suppose $B>0$ and there exists $\zeta\in\mathbb{R}^{+}$ such that $\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\zeta$. Let $\mathbf{F}\in\left\{\mathbf{CWTM},\mathbf{CWMed},\mathbf{GM},\mathbf{MK}\right\}$, $\gamma=\min\left\{\left(\frac{\Delta_{o}}{\kappa G^{2}}\right)\frac{1}{T},~{}\frac{1}{L}\right\}$ and $T\geq\frac{\Delta_{o}L}{\kappa G^{2}}$. Consider an arbitrary real value $\xi_{o}\in(0,1)$. Let $\rho\coloneqq\exp\left(\frac{(2+B^{2})\Delta_{o}}{(1-\xi_{o})G^{2}}L\right)\zeta$. Then, the following holds true:
  1. 1.
    If $\frac{f}{n}\coloneqq(1-\xi)\mathsf{BP}$, where $\xi\in(0,\xi_{o}]$, then, $\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}\left(\hat{{{\theta}}}\right)}\right\rVert^{2}}\right]\leq\xi\,\Psi(G,B,\rho)~{}\varepsilon_{o}$.
  2. 2.
    If $\mathsf{BP}\leq f/n<1/2$, then $\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}\left(\hat{{{\theta}}}\right)}\right\rVert^{2}}\right]\leq\Psi(G,B,\rho)\cdot\frac{G^{2}}{8}$
In order to prove Theorem C.4, we first obtain the following corollary of Lemma C.2 and Lemma 2.6.
Corollary C.5.
For the parameters given in Lemma C.2, if $T\geq\frac{\Delta_{o}L}{\kappa G^{2}}$, then Robust-DGD $\circ$ ARC achieves
$\mathop{{}\mathbb{E}}\left[{\left\lVert{\nabla\mathcal{L}_{\mathcal{H}}\left(\hat{{{\theta}}}\right)}\right\rVert^{2}}\right]\leq 5\kappa~{}\min\left\{G^{2}+B^{2}\rho^{2},~{}\frac{G^{2}}{\max\{(1-\kappa B^{2}),0\}}\right\},$
Proof of Corollary C.5.
For simplicity, we write $\mathcal{L}_{\mathcal{H}}$ as $\mathcal{L}$ throughout the proof.
Since $T\geq\frac{\Delta_{o}L}{\kappa G^{2}}$, $\gamma=\frac{\Delta_{o}}{\kappa G^{2}T}\leq\frac{1}{L}$. Therefore, from Lemma C.2 we obtain that
$\frac{1}{T}\sum_{t=1}^{T}\left\lVert{\nabla\mathcal{L}({{\theta}_{t}})}\right\rVert^{2}\leq 5\kappa\left(G^{2}+B^{2}\rho^{2}\right).$ (29)
@@
We are now ready to prove Theorem C.4. Recall that
$\mathsf{BP}\coloneqq\frac{1}{2+B^{2}},\quad\varepsilon_{o}\coloneqq\frac{1}{4}\cdot\frac{G^{2}(f/n)}{1-(2+B^{2})(f/n)},\quad\text{ and }$
@@
Proof of Theorem C.4.
For simplicity, we write $\mathcal{L}_{\mathcal{H}}$ as $\mathcal{L}$ throughout the proof.
@@
We first consider the case when $\frac{f}{n}=\frac{1-\xi}{2+B^{2}}$. In this particular case,
$n=\left(\frac{2}{1-\xi}+\frac{B^{2}}{1-\xi}\right)f\geq\left(2+\frac{B^{2}}{1-\xi}\right)f.$
Therefore, thanks to Lemma 2.5, $\mathbf{F}$ is $(f,\kappa)$-robust with
$\kappa\leq\frac{16f}{n-f}\left(1+\frac{1-\xi}{B^{2}}\right)^{2}=\frac{16(1-\xi)}{1+B^{2}+\xi}\left(1+\frac{1-\xi}{B^{2}}\right)^{2}\leq\frac{16(1-\xi)}{1+B^{2}}\left(1+\frac{1}{B^{2}}\right)^{2}\leq\kappa_{o}(1-\xi),$
@@
Summarizing from above, we have
@@
Therefore, since $\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\zeta$ for all $i\in\mathcal{H}$, the condition: $\max_{i\in\mathcal{H}}\left\lVert{\nabla\mathcal{L}_{i}({{\theta}_{1}})}\right\rVert\leq\exp\left(-\frac{\Delta_{o}}{\kappa G^{2}}L\right)\rho$ in Lemma C.2 is satisfied. Thus, by Corollary C.5 and (31) we obtain that
@@
Substituting from above in (32) concludes the proof for the first part of Theorem C.4.
The second part of Theorem C.4 follows immediately from the first inequality in (32), using the fact that $\kappa_{o}(1-\xi)\leq\kappa_{o}$. This concludes the proof. ∎
@@ Appendix D Comprehensive Experimental Setup
In this section, we present the comprehensive experimental setup considered in our paper.
@@ D.1 Datasets and Heterogeneity
In our experiments, we consider three standard image classification datasets, namely MNIST Deng (2012), Fashion-MNIST Xiao et al. (2017), and CIFAR-10 Krizhevsky et al. (2014). To simulate data heterogeneity in our experiments, we make the honest workers sample from the datasets using a Dirichlet Hsu et al. (2019a) distribution of parameter $\alpha$, as done in Hsu et al. (2019b); Allouah et al. (2023a); Farhadkhani et al. (2023). The smaller the $\alpha$, the more heterogeneous the setting. In our empirical evaluation, we set $\alpha\in\{0.1,0.5,1\}$ on (Fashion-)MNIST and $\alpha\in\{0.05,0.075,0.1,0.2,0.5\}$ on CIFAR-10 (refer to Figure 7). Furthermore, on MNIST and Fashion-MNIST, we also consider an extreme heterogeneity setting where the datapoints are sorted by increasing labels (0 to 9) and sequentially split equally among the honest workers.
@@ D.2 Algorithm, Distributed System, ML Models, and Hyperparameters
We perform our experiments using Robust-DSGD (see Algorithm 3), which is known to be an order-optimal robust variant of Robust-DGD Allouah et al. (2023a). On MNIST and Fashion-MNIST, we execute Robust-DSGD in a distributed system composed of $n-f=10$ honest workers, and $f\in\{1,2,3,4,5,7,9\}$ adversarial workers. Furthermore, we train a convolutional neural network (CNN) of 431,080 parameters with batch size $b=25$, $T=1000$, $\gamma=0.1$, and momentum parameter $\beta=0.9$. Moreover, the negative log likelihood (NLL) loss function is used, along with an $\ell_{2}$-regularization of $10^{-4}$. On CIFAR-10, we execute Robust-DSGD using a CNN of 1,310,922 parameters, in a distributed system comprising $n-f=16$ honest workers and $f=1$ adversarial worker. We set $b=50$, $T=2000$, $\beta=0.9$, and $\gamma=0.05$ decaying once at step 1500. Finally, we use the NLL loss function with an $\ell_{2}$ regularization of $10^{-2}$.
Algorithm 3 Robust Distributed Stochastic Gradient Descent (Robust-DSGD)
@@
In order to present the architectures of the ML models used in our experiments, we adopt the following compact terminology introduced in Allouah et al. (2023a).
@@ D.3 Byzantine Attacks
In our experiments, the adversarial workers execute five state-of-the-art adversarial attacks from the Byzantine ML literature, namely sign-flipping (SF) Allen-Zhu et al. (2020), label-flipping (LF) Allen-Zhu et al. (2020), mimic Karimireddy et al. (2022), fall of empires (FOE) Xie et al. (2020), and a little is enough (ALIE) Baruch et al. (2019).
@@
  • Mimic: the adversarial workers mimic a certain honest worker by sending its gradient to the server. In order to determine the optimal honest worker for the adversarial workers to mimic, we use the heuristic in Karimireddy et al. (2022).
@@ F.1 Limitations of Static Clipping
We perform our experiments using Robust-DSGD (see Algorithm 3) on (Fashion-)MNIST and CIFAR-10. We consider three different levels of heterogeneity, that we call moderate (corresponding to sampling from a Dirichlet distribution of parameter $\alpha=1$), high ($\alpha=0,1$), and extreme (as explained in Appendix D). On MNIST and Fashion-MNIST, we execute Robust-DSGD in a distributed system composed of $n=15$ workers, among which $f\in\{3,4\}$ are Byzantine. Furthermore, we train a convolutional neural network of 431,080 parameters with batch size $b=25$, $T=1000$, $\gamma=0.1$, and momentum parameter $\beta=0.9$. Moreover, the negative log likelihood loss function is used, along with an $\ell_{2}$-regularization of $10^{-4}$. On CIFAR-10, we execute Robust-DSGD on ResNet-18 He et al. (2015), in a distributed system comprising $n=9$ workers among which $f\in\{1,2\}$ are Byzantine. We set $b=128$, $T=2000$, $\beta=0.9$, and $\gamma=0.1$ decaying once by 10$\times$ at step 1500. Finally, we use the cross-entropy loss function with an $\ell_{2}$ regularization of $5\times 10^{-4}$.
@@
Unreliability due to the attack. First, the efficacy of static clipping is significantly influenced by the type of Byzantine attack executed during the learning process. To illustrate, while $C=2$ exhibits the best performance under the SF attack Allen-Zhu et al. (2020), depicted in Figure 1, it leads to a complete collapse of learning under LF Allen-Zhu et al. (2020). Consequently, identifying a static clipping threshold $C$ that consistently performs well in practice against any attack is difficult (if not impossible). This highlights the fragility of static clipping, as the unpredictable nature of the Byzantine attack, a parameter that cannot be a priori known, can significantly degrade the performance of the chosen static clipping approach.
@@ F.2 Empirical Benefits of ARC
ARC exhibits robust performance across diverse Byzantine scenarios, encompassing variations in both the type of Byzantine attack and the number of Byzantine workers $f$. As depicted in Figures 1 and 17, ARC consistently yields robust performance, regardless of the Byzantine attack. In extreme heterogeneity with $f=3$, ARC maintains an accuracy of 75%, while $C=2$ demonstrates a subpar worst-case accuracy of 25% (also refer to the right plot of Figure 1). Furthermore, despite the increase in the number of Byzantine workers to $f=4$, ARC remains the top-performing clipping approach among all considered strategies.